OpenStudy (anonymous):

Express in terms of i. sqrt of -2

OpenStudy (anonymous):

\[\sqrt{-2}=\sqrt{2i^{2}}=i \sqrt{2}\]

OpenStudy (anonymous):

add + or - before it